Broadband – 14:00 – 25/03/2021

We are seeing a number of broadband circuit’s dropping PPP and some intermittent packet loss. We are currently investigating.

UPDATE01 – 15:00

We have raised a case with our wholesale suppler and have moved traffic from Telehouse North to Telehouse East to see if this improves things.

UPDATE02 – 16:00

PPP reconnections have reduced but we are still seeing packet loss spikes to a large number of connections.

UPDATE03 – 18:00

We are still working with our wholesale supplier to understand the root cause of the issue as we have eliminated our network.

UPDATE04 – 22:00

We have chased our wholesale supplier for an update. Traffic remains via Telehouse East. The packet loss we are seeing does not seem to be service affecting at this stage but of course it should not be there.

UPDATE05 – 12:00 (26/03/2021)

We have dispatched an engineer to swap out optics on our side of the link.

UPDATE06 – 13:00

We have escalated this within wholesale as the problem remains. We do apologise for the inconvenience this is causing. We are still pending an optics swap as well.

UPDATE07 – 16:30

We have had a response from our wholesale suppler to advise additional resources have been added to our case. We are also still pending an optic swap from Telehouse.

UPDATE – 17:00

We have provided example cases of circuits where we are not seeing the same issue. We also believe this is only affecting a specific type of traffic.

UPDATE09 – 17:30

We have reached out to our hardware vendor to see if there is additional diagnostic tools can be provided. We apologise for the continued delay in getting this resolved.

UPDATE10 – 20:15

Optics have been changed at Telehouse North. Unfortunately the Telehouse engineer was not given an update from ourselves prior to starting the work which sadly resulted in traffic being dropped. We are continuing to monitor the interface.

UPDATE11 – 20:35 (29/03/2021)

We have observed some stability restore to the network over the past 72 hours. However we are still concerned there may be an issue on both London core devices caused by a Memory Leak and are working towards a maintenance window to eliminate this. Further details will be posted of the maintenance and times / impact as new “events / post” So they are clearly seen.

Broadband – 27/01/2021 – London 020

We are aware of a repeat Openreach fault from yesterday affecting broadband services within the 020 area code. We have raised this with our suppliers and awaiting an update.

UPDATE01 – 13:55

We have been advised the root cause has been found and a fix is being implemented.

UPDATE02 – 16:00

We have asked for an update from our supplier.

UPDATE03 – 17:10

We have been advised an update is not due before 19:00 We have gone back to advise this is unacceptable. Our service account manager at wholesale has stepped in to push Openreach for further details and a resolution time.

We apologize for the inconvenience this is causing.

UPDATE04 – 20:20

We have had updates provided from our wholesale supplier to advise while Openreach have not provided any firm ETA and raw fault detail. They believe the outage is being caused by an aggregated fibre node serving what they refer to as the Parent exchange and the secondary exchanges.

We are continuing to push for updates and proactively getting updates from wholesale now.

UPDATE05 – 02:30

We have been advised the fault has been resolved. We are awaiting an RFO and will publish once provided.

We apologise for the inconvenience.

04/01/2021 22:40 – Broadband Network

Our network monitoring has alerted us to one of our broadband LNS gateways reloading. This has resulted in broadband services connected via this device to disconnect and failover to alternative gateways.

The device has reloaded and returned to service. We are currently investigating the cause internally.

We apologize for any inconvenience this may have caused but do not expect a recurrence.

Network Maintenance 19/10/2020

We will be making some changes to our network on the 19/10/2020 starting from 12:00 noon. These changes are in relation to our LINX peering and no impact is expected to be seen as traffic will be gracefully re-routed via alternative routes prior to work starting.

We will also be taking this opportunity to add additional hardware for future broadband capacity. This will also be non service affecting.

UPDATE 01 – FINAL – 16:50

This work is complete and LINX peering has been restored. No impact was seen to customer traffic.

14/10/2020 12:43– Leased Lines

We are aware of an issue affecting leased lines services delivered via Telehouse North. We are investigating as a matter of urgency. We sorry for any inconvenience being caused.

Services with backup DSL will have been automaticity re-routed.

UPDATE 01 – 12:50

Further investigations have shown this is affecting services delivered from our redundant fibre routes as well. We are continuing to work with our suppliers as the root cause has been identified outside of our network.

UPDATE 02 – 13:02

We are starting to see services recover but services should still be considered at risk.

UPDATE 03 – 14:00

We have had further feedback from our suppliers to advise this has been resolved and they believe this to have been down to a configuration issue there side. We have raised this as concern as this took down multiple fail over links for redundancy.

FINAL – 10:39

We have been advise this issue was cause by human error within our wholesale provider and the failure to follow strict guidelines when undertaking work. Due to the impact this had on ourselves and the loss of our redundant links with this provider, We are undertaking an internal review to ensure we mitigate against this in the future.

14/09/2020 22:00 – Broadband Network

Due to a unexpected reload of a broadband gateway on our network this afternoon, we are seeing a traffic imbalance on part of our broadband network.

We will be taking steps to disconnect sessions gratefully and re balance the affected gateways.

End users will see a PPP reconnect taking approximately 5-20 seconds. In the rare event your connection does not restore. You will need to power OFF and ON your router.

UPDATE01 – 22:10

This work is now complete.

27/08/2020 – VoIP MSO [Resolved]

We are currently experiencing an issue with VoIP calls across our network, engineers are working on this and will provide an update asap.

-Update- 27/08/2020 09:48
The issue has been identified with one of our up stream carriers who are currently working on an issue and will provide updates shortly.

-Update- 27/08/2020 09:56 Telstra LHC Outage

All powered equipment in Telstra LHC Data Center went offline at 09:17 BST. We have been informed there are multiple fire engines on site and a suspected UPS fire on the 3rd floor where our comms equipment is located. It seems most likely the fire brigade have ordered building power down as part of electrical safety procedure.

As far as we are aware, this affects all customers and carriers within LHC, and we have confirmation that other carriers with passive connections extending outside the building are also showing offline. This is therefore affecting customers who are directly terminated with active circuits at this location. All passive fibre connections remain unaffected, including those passing through the building.

Updates to follow as they arrive from the DC. We sincerely apologise for any inconvenience this may cause.

-Update- 27/08/2020 11:10

Correspondence from LHC DC:
Due to a localised fire in the LHC, we have lost the Green system. This provides power to the lower half of the building. The fire has tripped the breakers supporting the busbar. Engineers are on-site and are working to restore power via generator as we speak.

-Update- 27/08/2020 13:16

We have been made aware by the DC that the services are starting to restore. We are monitoring this carefully and will provide you with an update as soon as we have more information.

-Update- 27/08/2020 14:08

We are seeing call flows return to normal levels but have yet to hear back from the DC and/or Carrier. We will continue to monitor and provide updates as they become available.

-Update- 27/08/2020 16:07

Engineers are still working to improve stability in the network and restore the remaining services we hope this will be complete with in the next 30 minutes. Most customers should have service restored.

-Update- 27/08/2020 16:40

We can see normal call flow across the network but have still yet to get a clear message from the carrier.

-Update- 27/08/2020 17:45

The carrier reports all services are now restored.

-Update- 28/08/2020 9:30

Services are now fully restored, a full RFO will be posted once made available.

London Data Center Fire – 07:25 – 18/08/2020

We have seen a number of leased lines and peering sessions go down over night. This has been caused by a possible fire (fire alarms are going off) in at least 1 London data center affecting a number of providers. We are working to obtain further information.

UPDATE 01 – 8:00

We have been advised the London harbor building (Equinix LD8)remains evacuated.

UPDATE 02 – 09:15

Equinix have advised that a fire alarm was triggered by the failure of output static switch from their Galaxy UPS system. This has resulted in a loss of power for multiple customers and Equinix IBX Engineers are working to resolve the issue and restore power. At this moment in time we do not believe there to have been a fire.

UPDATE 03 – 10:15

Equinix IBX Site Staff report that the root cause of the fire evacuation was due to the failure of a Galaxy UPS that triggered the fire alarm. The fire system has been reinstated and IBX Staff have been allowed back in to the building. We are now awaiting updates on restoring services.

UPDATE 04 – 11:15

Equinix Engineers have advised that their IBX team have begun restoring power to affected devices. Unfortunately, at present there remains no estimate resolution time.

UPDATE 05 – 12:15

Equinix have advised that services are started to be restored with equipment being migrated over to other newly installed infrastructure. We have yet to see any of our affected connections restore but keep checking for updates.

UPDATE 06 – 13:15

Equinix IBX Site Staff reports that services have been further restored to more customers and IBX Engineers continue to work towards restoring services to all customers by migrating to the newly installed and commissioned infrastructure. Equinix advised access will be granted and prioritized to the IBX should any customers need it to work on their equipment.

UPDATE 07 – 14:20

IBX Site Staff reports that services have been further restored to more customers and increasing numbers of those affected are now operational along with the majority of Equinix Network Services. IBX Engineers continue to work towards restoring services to all customers by migrating to the newly installed and commissioned infrastructure

UPDATE 08 – 15:15

We are please to advise we have just seen all affected services restore. Circuits remain at risk due to the ongoing power issues on site, however we do not expect them to go down again.